# Quantum-private distributed sensing

### v1

@@ 5 Analysis of privacy and integrity
$\varepsilon_{p}\leq\frac{24}{n^{2}}\sqrt{\frac{2\sqrt{c}}{n}+2nf},$ (SM1)
where $n$ is the number of sensors and $f$ is the failure rate of the stabilizer tests, averaged over all the $n$ tested stabilizers; conditional on the success from the verification protocol we note that $f\leq 1/(2n^{2})$. The positive constant $c$ is a free variable constrained by $43/2m<c<(n-1)^{2}/$ 4. The verifier can choose the failure rate $f$ and the free variable $c$ to obtain a satisfactory level of privacy and integrity in their protocol and decrease $\varepsilon_{p}$ to an arbitrarily small value.
In particular, $\varepsilon_{p}$-privacy quantifies the local sensor information leakage to an adversary. We extend the privacy definition from [25] to the case of $n$ parameters encoded on a multipartite state via unitary encoding.
@@
$\quad F_{Q|\theta_{j}}(\rho_{\phi_{1},\dots,\theta_{j},\dots,\phi_{n}})\leq\varepsilon_{p}(h_{\max}-h_{\min})^{2},$ (SM2)
@@
$\quad F_{Q|\theta_{j}}(\rho_{\phi_{1},\dots,\theta_{j},\dots,\phi_{n}})\leq\varepsilon_{p}n^{2}.$ (SM3)
@@
We now discuss the $\varepsilon_{i}$-integrity of the protocol. The authors of Ref. [25] quantify the deviation in accuracy and in precision of the estimator of the protocol, when the shared resource deviates from a GHZ state. As such, these measures depend on an upper bound on the average trace distance between the encoded resource state and the encoded GHZ state –averaged over the parameter estimation rounds [25]. Thus, we have the following definition.
@@
$\tau_{\phi}=\ket{\Phi}\bra{\Phi},\quad\ket{\Phi}=\frac{1}{\sqrt{2}}(|0\rangle^{\otimes n}+e^{i\phi}|1\rangle^{\otimes n}).$ (SM5)
The deviations in accuracy and precision derived in [25] are obtained for a specific estimator $\hat{\phi}$ of the global phase. In particular, let $\tau_{\phi}$ be the ideal encoded state, (SM5). Then, the expectation value of $\sigma_{X}^{\otimes n}$ on $\tau_{\phi}$ is:
$g(\phi)=\text{Tr}[\sigma_{X}^{\otimes n}\tau_{\phi}]=\cos\phi.$ (SM6)
@@
$=\arccos(\hat{g}).$ (SM8)
The authors in [25] argue that the estimator in (SM8), for $\nu\gg 1$, is unbiased when the resource state coincides with the ideal state ($\tau_{j}=\tau_{\phi}$ for all $j$). Then, Theorem 1 in [25] provides a bound on the bias of the estimator computed on the actual resource state:
$|\mathbb{E}(\hat{\phi})-\phi|\leq\frac{2o\varepsilon_{i}}{|\sin\phi|},$ (SM9)
where $\mathbb{E}(\hat{\phi})$ is the expectation value of (SM8) on the actual resource state, while $o$ is the maximum magnitude of the eigenvalues of the observable which is being measured for parameter estimation; in our case $o=1$.
Similarly, one can argue that the variance of the estimator in (SM8), assuming that the resource state is the same in every parameter estimation round ($\tau_{j}=\tau$), can be computed as [25]:
$\Delta^{2}\hat{\phi}=\frac{1-(\text{Tr}[\sigma_{X}^{\otimes n}\tau])^{2}}{\nu\,\sin^{2}\phi}.$ (SM10)
@@
$\left\lvert\Delta^{2}\hat{\phi}-\frac{1}{\nu}\right\rvert\leq\frac{4\varepsilon_{i}(2\nu^{-1}+\varepsilon_{i})}{\sin^{2}\phi}.$ (SM11)
We evaluate the bounds in (SM11) and (SM9) by replacing $\mathbb{E}(\hat{\phi})$ with $\hat{\phi}$ and $\Delta^{2}\hat{\phi}$ with the formula in (SM10), where we choose $\tau=\sum_{j}\tau_{j}/\nu$:
$\Delta^{2}\hat{\phi}$ $=\frac{1-(\mathbb{E}(\hat{g}))^{2}}{\nu\,\sin^{2}\phi}$ (SM12)
$\approx\frac{1-(\hat{g})^{2}}{\nu\,\sin^{2}\phi}.$ (SM13)
@@
Figure SM1: The accuracy and precision of the protocol, defined with respect to $\varepsilon_{i}$. In both plots, the green line is defined w.r.t $\varepsilon_{i,\textrm{theo}}=\sqrt{2\sqrt{c}/n+2nf}$ and the purple line is defined w.r.t $\varepsilon_{i,\textrm{exp}}=\sqrt{1-F(\rho,\sigma)}$. (a) shows the accuracy, as in (SM9), of the measured global phase (black line). Specifically, the green and purple are the right hand side of the inequality and the black line is the left hand side of the inequality in (SM9). (b) shows the precision, a more direct measure of integrity as in (SM11) it is w.r.t a certain number of parameter estimation rounds $\nu$. The green and purple lines are the right hand side and the black line is the left hand side of the inequality in (SM11). This is taken at $\nu=3200$, $c=0.25$, with $f=0.047$ as measured in (SM9), and the fidelity used is 0.905, the lower bound of 0.907(2).
@@ 7 Direct calculation of QFI for privacy statement
In this section we perform a direct calculation of the QFI appearing in the privacy definition that applies to our experiment, Definition 2. This is achieved by employing the tomographically-reconstructed resource state of our experiment. The goal is to compare the direct calculation of the QFI, which would lead to a tight privacy bound $\varepsilon_{p}$, with the upper bound on $\varepsilon_{p}$ provided by [25] and reported in (SM1).
@@
$F_{Q|\theta}(\rho_{\theta})=2\sum_{k,l}\frac{\left\lvert\braket{\psi_{k}}{\partial_{\theta}\rho_{\theta}}{\psi_{l}}\right\rvert^{2}}{\lambda_{k}+\lambda_{l}},$ (SM14)
@@
$\rho_{\theta}=e^{-i\theta A}\rho e^{i\theta A},$ (SM15)
@@
$F_{Q|\theta}(\rho_{\theta})=2\sum_{k,l}\frac{(\lambda_{k}-\lambda_{l})^{2}}{\lambda_{k}+\lambda_{l}}\left\lvert\braket{\varphi_{k}}{A}{\varphi_{l}}\right\rvert^{2},$ (SM16)
where now $\lambda_{k}$ and |$\ket{\varphi_{k}}$ are obtained from the spectral decomposition of the unencoded state $\rho$: $\rho=\sum_{k}\lambda_{k}\ket{\varphi_{k}}\bra{\varphi_{k}}$ |, implying that the QFI in (SM16) is independent of the value of $\theta$. Moreover, if the state $\rho$ is pure, $\rho=\ket{\varphi}\bra{\varphi}$ |, it holds:
$F_{Q|\theta}(\rho_{\theta})=4\left[\braket{\varphi}{A^{2}}{\varphi}-\left(\braket{\varphi}{A}{\varphi}\right)^{2}\right].$ (SM17)
In order to determine the privacy parameter $\varepsilon_{p}$, let us fix for the moment $j=1$ in (SM3), implying that we want to make a statement about the privacy of the local phase $\theta_{1}$ when encoded in our resource state $\rho$. In order to find the best possible privacy parameter $\varepsilon_{1}$, we solve the following optimization problem:
$\alpha_{1}=\min_{\phi_{k}(\theta),\,k\neq 1}F_{Q|\theta_{1}}(\rho_{\theta_{1},\phi_{2}(\theta_{1}),\dots,\phi_{n}(\theta_{1})}),$ (SM18)
and then choose $\varepsilon_{1}=\alpha_{1}/n^{2}$. The optimization is done over all possible choices of functions $\phi_{k}(\theta)$, such that the resulting state is functionally independent of $\theta_{1}$ –or as independent as possible. If we repeat this procedure for each $j$, we can then set $\varepsilon=\max_{j}\varepsilon_{j}$ and claim that the resource state $\rho$ is $\varepsilon_{p}$-private according to definition (SM3). To solve the optimization in (SM18), we compute the QFI, with respect to $\theta_{1}$, of the encoded resource state:
$\rho_{\theta_{1},\phi_{2}(\theta_{1}),\dots,\phi_{n}(\theta_{1})}=e^{-i\theta_{1}\frac{Z_{1}}{2}}e^{-i\phi_{2}(\theta_{1})\frac{Z_{2}}{2}}\cdots e^{-i\phi_{n}(\theta_{1})\frac{Z_{n}}{2}}\,\rho\,e^{i\theta_{1}\frac{Z_{1}}{2}}e^{i\phi_{2}(\theta_{1})\frac{Z_{2}}{2}}\cdots e^{i\phi_{n}(\theta_{1})\frac{Z_{n}}{2}},$ (SM19)
where $\rho$ is the resource state outputted by the verification procedure. To this aim, we start from the QFI formula in (SM14) and derive a generalized version of (SM16) which is valid when the state is encoded with multiple parameters via unitary encoding:
$\rho_{\theta}=e^{-i\theta A_{1}}e^{-i\phi_{2}(\theta)A_{2}}\cdots e^{-i\phi_{n}(\theta)A_{n}}\,\rho\,e^{i\theta A_{1}}e^{i\phi_{2}(\theta)A_{2}}\cdots e^{i\phi_{n}(\theta)A_{n}},$ (SM20)
with the additional assumption that the generators $A_{j}$ commute pairwise. Then, the encoded resource state in (SM19) is a particular case of (SM20). We start by computing the derivative of the state in (SM20) with respect to $\theta$:
$\partial_{\theta}\rho_{\theta}=U_{\theta}[-iA_{1},\rho]U^{\dagger}_{\theta}+\sum_{k=2}^{n}U_{\theta}[-iA_{k}\phi^{\prime}_{k},\rho]U^{\dagger}_{\theta},$ (SM21)
@@
$U_{\theta}:=e^{-i\theta A_{1}}e^{-i\phi_{2}(\theta)A_{2}}\cdots e^{-i\phi_{n}(\theta)A_{n}}.$ (SM22)
@@
$=i(\lambda_{m}-\lambda_{l})\left(\braket{\varphi_{m}}{A_{1}}{\varphi_{l}}+\sum_{k=2}^{n}\phi^{\prime}_{k}(\theta)\braket{\varphi_{m}}{A_{k}}{\varphi_{l}}\right),$ (SM23)
which substituted in (SM14) yields:
$F_{Q|\theta}(\rho_{\theta})=2\sum_{m,l}\frac{(\lambda_{m}-\lambda_{l})^{2}}{\lambda_{m}+\lambda_{l}}\left\lvert\braket{\varphi_{m}}{A_{1}}{\varphi_{l}}+\sum_{k=2}^{n}\phi^{\prime}_{k}(\theta)\braket{\varphi_{m}}{A_{j}}{\varphi_{l}}\right\rvert^{2},$ (SM24)
which is the QFI of the state (SM20), with commuting generators and where $\lambda_{l}$ and |$\ket{\varphi_{l}}$ are obtained from the spectral decomposition of the unencoded state: $\rho=\sum_{l}\lambda_{l}\ket{\varphi_{l}}\bra{\varphi_{l}}$ |. Now, we apply the last expression for the state in (SM19) to solve the optimization problem in (SM18). We obtain:
$\alpha_{1}=\frac{1}{2}\min_{\phi_{k}(\theta),\,k\neq 1}\sum_{m,l}\frac{(\lambda_{m}-\lambda_{l})^{2}}{\lambda_{m}+\lambda_{l}}\left\lvert\braket{\varphi_{m}}{Z_{1}}{\varphi_{l}}+\sum_{k=2}^{n}\phi^{\prime}_{k}(\theta)\braket{\varphi_{m}}{Z_{k}}{\varphi_{l}}\right\rvert^{2},$ (SM25)
where the solution of the optimization, $\alpha_{1}$, can potentially depend on $\theta$ if the derivatives of the functions $\phi_{k}(\theta)$ are not constant. However, we can assume the functions $\phi_{k}(\theta)$ to be of the form: $\phi_{k}(\theta)=c_{k}\theta$, for $c_{k}\in\mathbbm{R}$, without loss of generality. Indeed, suppose that $\tilde{\phi}_{k}(\theta)$ are the optimal functions that solve the optimization problem, and suppose that $\alpha_{1}(\theta)$ depends on $\theta$ through $\tilde{\phi}^{\prime}_{k}(\theta)$. Let $\bar{\theta}$ be the parameter that minimizes $\alpha_{1}$, i.e.: $\bar{\theta}=\arg\min_{\theta}\alpha_{1}(\theta)$. Then, we can choose another set of functions directly proportional to $\theta$, namely $\phi_{k}(\theta)=\theta\tilde{\phi}^{\prime}_{k}(\bar{\theta})$, such that the resulting QFI is minimal and equal to $\alpha_{1}(\bar{\theta})$, with the added benefit of being independent of $\theta$. Therefore, we can restrict the optimization in (SM25) to linear functions of $\theta$. Thus, we get the following parameter for the privacy of $\theta_{1}$ when encoded in the resource state:
$\varepsilon_{1}=\frac{1}{2n^{2}}\min_{c_{k}\in\mathbbm{R}}\sum_{m,l}\frac{(\lambda_{m}-\lambda_{l})^{2}}{\lambda_{m}+\lambda_{l}}\left\lvert\braket{\varphi_{m}}{Z_{1}}{\varphi_{l}}+\sum_{k=2}^{n}c_{k}\braket{\varphi_{m}}{Z_{k}}{\varphi_{l}}\right\rvert^{2}.$ (SM26)
@@
$\varepsilon_{p}=\frac{1}{2n^{2}}\max_{1\leq j\leq n}\,\,\min_{c_{k|j}\in\mathbbm{R}}\sum_{m,l}\frac{(\lambda_{m}-\lambda_{l})^{2}}{\lambda_{m}+\lambda_{l}}\left\lvert\braket{\varphi_{m}}{Z_{j}}{\varphi_{l}}+\sum_{k\neq j}c_{k|j}\braket{\varphi_{m}}{Z_{k}}{\varphi_{l}}\right\rvert^{2}.$ (SM27)
@@
$\rho_{\phi_{1},\dots,\theta_{j},\dots,\phi_{n}}=U_{1}U_{2}\cdots U_{n}\ket{GHZ}\bra{GHZ}U^{\dagger}_{1}U^{\dagger}_{2}\cdots U^{\dagger}_{n}=\ket{GHZ}\bra{GHZ}.$ | . (SM28)
Thus, by definition of the QFI in (SM14), it holds: $F_{Q|\theta_{j}}(\rho_{\phi_{1},\dots,\theta_{j},\dots,\phi_{n}})=0$ for every $j$ and hence $\varepsilon_{p}=0$ according to Definition 1. Alternatively, we can directly evaluate the privacy parameter from (SM27). For this, we start from the spectral decomposition of the GHZ state (where we identified |$\ket{\varphi_{1}}=\ket{GHZ}$):
@@
which employed in (SM27) yields:
@@
$=\frac{1}{n^{2}}\max_{1\leq j\leq n}\,\,\min_{c_{k|j}\in\mathbbm{R}}\left[\braket{GHZ}{(Z^{\prime}_{j})^{2}}{GHZ}-\left(\braket{GHZ}{Z^{\prime}_{j}}{GHZ}\right)^{2}\right],$ (SM30)

### v2

@@ 5 Analysis of privacy and integrity
In the main body, we introduced equation 1 which related the lower-bound fidelity of the output state $\rho_{j}$ to the GHZ state when performing the Verification protocol with parameters $n,f,c,m$. One can rewrite this expression to relate this lower-bounded fidelity to the quantum Fisher information by employing the continuity of QFI as shown in Ref. [27] to arrive at,
$\mathbb{P}\bigg{(}F_{Q|\theta}(\rho_{\theta})\leq 24\frac{2\sqrt{c}}{n}+2nf\Bigg{)}\geq 1-n^{1-\frac{2mc}{3}}.$ (SM1)
Where $F_{Q|\theta}(\rho_{\theta})$ is the QFI of parameter $\theta$ given the state $\rho_{\theta}$ while the remaining parameters are the same as in equation 1 of the main body. We remark that this expression corresponds to Theorem 2 in Ref. [27], although there was a sign error on the term $2nf$. It was subsequently shown that the QFI of the local parameters $\theta_{i}$ is zero for the GHZ state. However, this is only true in the ideal setting, instead one can bound the amount of privacy with the parameter $\varepsilon_{p}$ as $F_{Q|\theta}(\rho_{\theta})\leq\varepsilon_{p}n^{2}$, note that $n^{2}$ is used to reflect the maximum achievable QFI for an $n$-partite GHZ state. In this section, we will outline how the parameters $n,f,c,m$ impacts the privacy regarding the network sensing task as well as the accuracy for estimating the global phase when using the upper-bounded privacy parameter $\varepsilon_{p}$.
@@
$\varepsilon_{p}\leq\frac{24}{n^{2}}\sqrt{\frac{2\sqrt{c}}{n}+2nf},$ (SM2)
where $n$ is the number of sensors and $f$ is the failure rate of the stabilizer tests, averaged over all the $n+1$ stabilizers; conditional on the success from the Verification protocol which must satisfy $f\leq 1/(2n^{2})$. The positive constant $c$ is a variable constrained by $43/(2m)<c<(n-1)^{2}/$ 4, which should ideally be minimised to decrease the upper-bounded $\varepsilon_{p}$ value. Notably, the lower bound on $c$ is set by $m$ which directly relates to the number of test copies used, as captured by $N_{t}=\lceil 2mn^{5}\text{log}(n)\rceil$. In the main body we assumed infinite copies ($N_{t}\rightarrow\infty$), thus $c\rightarrow 0$, which simplifies equation SM2 to $\varepsilon_{p}\leq 24n^{-2}\sqrt{2nf}$. The verifier can then choose the failure rate $f$ to accept in the Verification protocol to obtain a satisfactory level of privacy and integrity in their protocol as this decreases $\varepsilon_{p}$.
In particular, $\varepsilon_{p}$-privacy quantifies the local sensor information leakage to an adversary. We extend the privacy definition from Ref. [27] to the case of $n$ parameters encoded on a multi-partite state via unitary encoding.
@@
$\quad F_{Q|\theta_{j}}(\rho_{\phi_{1},\dots,\theta_{j},\dots,\phi_{n}})\leq\varepsilon_{p}(h_{\max}-h_{\min})^{2},$ (SM3)
@@
$\quad F_{Q|\theta_{j}}(\rho_{\phi_{1},\dots,\theta_{j},\dots,\phi_{n}})\leq\varepsilon_{p}n^{2}.$ (SM4)
@@
We now discuss the $\varepsilon_{i}$-integrity of the protocol. The authors of Ref. [27] quantify the deviation in accuracy and in precision of the estimator of the protocol, when the shared resource deviates from a GHZ state. As such, these measures depend on an upper bound on the average trace distance between the encoded resource state and the encoded GHZ state—averaged over the parameter estimation rounds [27]. Thus, we have the following definition.
@@
$\tau_{\phi}=\ket{\Phi}\bra{\Phi},\quad\ket{\Phi}=\frac{1}{\sqrt{2}}(|0\rangle^{\otimes n}+e^{i\phi}|1\rangle^{\otimes n}).$ (SM6)
The deviations in accuracy and precision derived in [27] are obtained for a specific estimator $\hat{\phi}$ of the global phase. In particular, let $\tau_{\phi}$ be the ideal encoded state, (SM6). Then, the expectation value of $\sigma_{X}^{\otimes n}$ on $\tau_{\phi}$ is:
$g(\phi)=\text{Tr}[\sigma_{X}^{\otimes n}\tau_{\phi}]=\cos\phi.$ (SM7)
@@
$=\arccos(\hat{g}).$ (SM9)
The authors in [27] argue that the estimator in (SM9), for $\nu\gg 1$, is unbiased when the resource state coincides with the ideal state ($\tau_{j}=\tau_{\phi}$ for all $j$). Then, Theorem 1 in [27] provides a bound on the bias of the estimator computed on the actual resource state:
$|\mathbb{E}(\hat{\phi})-\phi|\leq\frac{2o\varepsilon_{i}}{|\sin\phi|},$ (SM10)
where $\mathbb{E}(\hat{\phi})$ is the expectation value of (SM9) on the actual resource state, while $o$ is the maximum magnitude of the eigenvalues of the observable which is being measured for parameter estimation; in our case $o=1$.
Similarly, one can argue that the variance of the estimator in (SM9), assuming that the resource state is the same in every parameter estimation round ($\tau_{j}=\tau$), can be computed as [27]:
$\Delta^{2}\hat{\phi}=\frac{1-(\text{Tr}[\sigma_{X}^{\otimes n}\tau])^{2}}{\nu\,\sin^{2}\phi}.$ (SM11)
@@
$\left\lvert\Delta^{2}\hat{\phi}-\frac{1}{\nu}\right\rvert\leq\frac{4\varepsilon_{i}(2\nu^{-1}+\varepsilon_{i})}{\sin^{2}\phi}.$ (SM12)
We evaluate the bounds in (SM12) and (SM10) by replacing $\mathbb{E}(\hat{\phi})$ with $\hat{\phi}$ and $\Delta^{2}\hat{\phi}$ with the formula in (SM11), where we choose $\tau=\sum_{j}\tau_{j}/\nu$:
$\Delta^{2}\hat{\phi}$ $=\frac{1-(\mathbb{E}(\hat{g}))^{2}}{\nu\,\sin^{2}\phi}$ (SM13)
$\approx\frac{1-(\hat{g})^{2}}{\nu\,\sin^{2}\phi}.$ (SM14)
@@
Figure SM1: The accuracy and precision of the protocol, defined with respect to $\varepsilon_{i}$. In both plots, the green line is defined w.r.t $\varepsilon_{i,\textrm{theo}}=\sqrt{2\sqrt{c}/n+2nf}$ and the purple line is defined w.r.t $\varepsilon_{i,\textrm{exp}}=\sqrt{1-F(\rho,\sigma)}$. (a) shows the accuracy, as in (SM10), of the measured global phase (black line). Specifically, the green and purple are the right hand side of the inequality and the black line is the left hand side of the inequality in (SM10). (b) shows the precision, a more direct measure of integrity as in (SM12) it is w.r.t a certain number of parameter estimation rounds $\nu$. The green and purple lines are the right hand side and the black line is the left hand side of the inequality in (SM12). This is taken at $\nu=3200$, $c=0.25$, with $f=0.047$ as measured in (SM10), and the fidelity used is 0.905, the lower bound of 0.907(2).
@@ 7 Direct calculation of QFI for privacy statement
In this section we perform a direct calculation of the QFI appearing in the privacy definition that applies to our experiment, Definition 2. This is achieved by employing the tomographically-reconstructed resource state of our experiment. The goal is to compare the direct calculation of the QFI, which would lead to a tight privacy bound $\varepsilon_{p}$, with the upper bound on $\varepsilon_{p}$ provided by [27] and reported in (SM2).
@@
$F_{Q|\theta}(\rho_{\theta})=2\sum_{k,l}\frac{\left\lvert\braket{\psi_{k}}{\partial_{\theta}\rho_{\theta}}{\psi_{l}}\right\rvert^{2}}{\lambda_{k}+\lambda_{l}},$ (SM15)
@@
$\rho_{\theta}=e^{-i\theta A}\rho e^{i\theta A},$ (SM16)
@@
$F_{Q|\theta}(\rho_{\theta})=2\sum_{k,l}\frac{(\lambda_{k}-\lambda_{l})^{2}}{\lambda_{k}+\lambda_{l}}\left\lvert\braket{\varphi_{k}}{A}{\varphi_{l}}\right\rvert^{2},$ (SM17)
where now $\lambda_{k}$ and |$\ket{\varphi_{k}}$ are obtained from the spectral decomposition of the unencoded state $\rho$: $\rho=\sum_{k}\lambda_{k}\ket{\varphi_{k}}\bra{\varphi_{k}}$ |, implying that the QFI in (SM17) is independent of the value of $\theta$. Moreover, if the state $\rho$ is pure, $\rho=\ket{\varphi}\bra{\varphi}$ |, it holds:
$F_{Q|\theta}(\rho_{\theta})=4\left[\braket{\varphi}{A^{2}}{\varphi}-\left(\braket{\varphi}{A}{\varphi}\right)^{2}\right].$ (SM18)
In order to determine the privacy parameter $\varepsilon_{p}$, let us fix for the moment $j=1$ in (SM4), implying that we want to make a statement about the privacy of the local phase $\theta_{1}$ when encoded in our resource state $\rho$. In order to find the best possible privacy parameter $\varepsilon_{1}$, we solve the following optimization problem:
$\alpha_{1}=\min_{\phi_{k}(\theta),\,k\neq 1}F_{Q|\theta_{1}}(\rho_{\theta_{1},\phi_{2}(\theta_{1}),\dots,\phi_{n}(\theta_{1})}),$ (SM19)
and then choose $\varepsilon_{1}=\alpha_{1}/n^{2}$. The optimization is done over all possible choices of functions $\phi_{k}(\theta)$, such that the resulting state is functionally independent of $\theta_{1}$ –or as independent as possible. If we repeat this procedure for each $j$, we can then set $\varepsilon=\max_{j}\varepsilon_{j}$ and claim that the resource state $\rho$ is $\varepsilon_{p}$-private according to definition (SM4). To solve the optimization in (SM19), we compute the QFI, with respect to $\theta_{1}$, of the encoded resource state:
$\rho_{\theta_{1},\phi_{2}(\theta_{1}),\dots,\phi_{n}(\theta_{1})}=e^{-i\theta_{1}\frac{Z_{1}}{2}}e^{-i\phi_{2}(\theta_{1})\frac{Z_{2}}{2}}\cdots e^{-i\phi_{n}(\theta_{1})\frac{Z_{n}}{2}}\,\rho\,e^{i\theta_{1}\frac{Z_{1}}{2}}e^{i\phi_{2}(\theta_{1})\frac{Z_{2}}{2}}\cdots e^{i\phi_{n}(\theta_{1})\frac{Z_{n}}{2}},$ (SM20)
where $\rho$ is the resource state outputted by the verification procedure. To this aim, we start from the QFI formula in (SM15) and derive a generalized version of (SM17) which is valid when the state is encoded with multiple parameters via unitary encoding:
$\rho_{\theta}=e^{-i\theta A_{1}}e^{-i\phi_{2}(\theta)A_{2}}\cdots e^{-i\phi_{n}(\theta)A_{n}}\,\rho\,e^{i\theta A_{1}}e^{i\phi_{2}(\theta)A_{2}}\cdots e^{i\phi_{n}(\theta)A_{n}},$ (SM21)
with the additional assumption that the generators $A_{j}$ commute pairwise. Then, the encoded resource state in (SM20) is a particular case of (SM21). We start by computing the derivative of the state in (SM21) with respect to $\theta$:
$\partial_{\theta}\rho_{\theta}=U_{\theta}[-iA_{1},\rho]U^{\dagger}_{\theta}+\sum_{k=2}^{n}U_{\theta}[-iA_{k}\phi^{\prime}_{k},\rho]U^{\dagger}_{\theta},$ (SM22)
@@
$U_{\theta}:=e^{-i\theta A_{1}}e^{-i\phi_{2}(\theta)A_{2}}\cdots e^{-i\phi_{n}(\theta)A_{n}}.$ (SM23)
@@
$=i(\lambda_{m}-\lambda_{l})\left(\braket{\varphi_{m}}{A_{1}}{\varphi_{l}}+\sum_{k=2}^{n}\phi^{\prime}_{k}(\theta)\braket{\varphi_{m}}{A_{k}}{\varphi_{l}}\right),$ (SM24)
which substituted in (SM15) yields:
$F_{Q|\theta}(\rho_{\theta})=2\sum_{m,l}\frac{(\lambda_{m}-\lambda_{l})^{2}}{\lambda_{m}+\lambda_{l}}\left\lvert\braket{\varphi_{m}}{A_{1}}{\varphi_{l}}+\sum_{k=2}^{n}\phi^{\prime}_{k}(\theta)\braket{\varphi_{m}}{A_{j}}{\varphi_{l}}\right\rvert^{2},$ (SM25)
which is the QFI of the state (SM21), with commuting generators and where $\lambda_{l}$ and |$\ket{\varphi_{l}}$ are obtained from the spectral decomposition of the unencoded state: $\rho=\sum_{l}\lambda_{l}\ket{\varphi_{l}}\bra{\varphi_{l}}$ |. Now, we apply the last expression for the state in (SM20) to solve the optimization problem in (SM19). We obtain:
$\alpha_{1}=\frac{1}{2}\min_{\phi_{k}(\theta),\,k\neq 1}\sum_{m,l}\frac{(\lambda_{m}-\lambda_{l})^{2}}{\lambda_{m}+\lambda_{l}}\left\lvert\braket{\varphi_{m}}{Z_{1}}{\varphi_{l}}+\sum_{k=2}^{n}\phi^{\prime}_{k}(\theta)\braket{\varphi_{m}}{Z_{k}}{\varphi_{l}}\right\rvert^{2},$ (SM26)
where the solution of the optimization, $\alpha_{1}$, can potentially depend on $\theta$ if the derivatives of the functions $\phi_{k}(\theta)$ are not constant. However, we can assume the functions $\phi_{k}(\theta)$ to be of the form: $\phi_{k}(\theta)=c_{k}\theta$, for $c_{k}\in\mathbbm{R}$, without loss of generality. Indeed, suppose that $\tilde{\phi}_{k}(\theta)$ are the optimal functions that solve the optimization problem, and suppose that $\alpha_{1}(\theta)$ depends on $\theta$ through $\tilde{\phi}^{\prime}_{k}(\theta)$. Let $\bar{\theta}$ be the parameter that minimizes $\alpha_{1}$, i.e.: $\bar{\theta}=\arg\min_{\theta}\alpha_{1}(\theta)$. Then, we can choose another set of functions directly proportional to $\theta$, namely $\phi_{k}(\theta)=\theta\tilde{\phi}^{\prime}_{k}(\bar{\theta})$, such that the resulting QFI is minimal and equal to $\alpha_{1}(\bar{\theta})$, with the added benefit of being independent of $\theta$. Therefore, we can restrict the optimization in (SM26) to linear functions of $\theta$. Thus, we get the following parameter for the privacy of $\theta_{1}$ when encoded in the resource state:
$\varepsilon_{1}=\frac{1}{2n^{2}}\min_{c_{k}\in\mathbbm{R}}\sum_{m,l}\frac{(\lambda_{m}-\lambda_{l})^{2}}{\lambda_{m}+\lambda_{l}}\left\lvert\braket{\varphi_{m}}{Z_{1}}{\varphi_{l}}+\sum_{k=2}^{n}c_{k}\braket{\varphi_{m}}{Z_{k}}{\varphi_{l}}\right\rvert^{2}.$ (SM27)
@@
$\varepsilon_{p}=\frac{1}{2n^{2}}\max_{1\leq j\leq n}\,\,\min_{c_{k|j}\in\mathbbm{R}}\sum_{m,l}\frac{(\lambda_{m}-\lambda_{l})^{2}}{\lambda_{m}+\lambda_{l}}\left\lvert\braket{\varphi_{m}}{Z_{j}}{\varphi_{l}}+\sum_{k\neq j}c_{k|j}\braket{\varphi_{m}}{Z_{k}}{\varphi_{l}}\right\rvert^{2}.$ (SM28)
@@
$\rho_{\phi_{1},\dots,\theta_{j},\dots,\phi_{n}}=U_{1}U_{2}\cdots U_{n}\ket{GHZ}\bra{GHZ}U^{\dagger}_{1}U^{\dagger}_{2}\cdots U^{\dagger}_{n}=\ket{GHZ}\bra{GHZ}.$ | . (SM29)
Thus, by definition of the QFI in (SM15), it holds: $F_{Q|\theta_{j}}(\rho_{\phi_{1},\dots,\theta_{j},\dots,\phi_{n}})=0$ for every $j$ and hence $\varepsilon_{p}=0$ according to Definition 1. Alternatively, we can directly evaluate the privacy parameter from (SM28). For this, we start from the spectral decomposition of the GHZ state (where we identified |$\ket{\varphi_{1}}=\ket{GHZ}$):
@@
which employed in (SM28) yields:
@@
$=\frac{1}{n^{2}}\max_{1\leq j\leq n}\,\,\min_{c_{k|j}\in\mathbbm{R}}\left[\braket{GHZ}{(Z^{\prime}_{j})^{2}}{GHZ}-\left(\braket{GHZ}{Z^{\prime}_{j}}{GHZ}\right)^{2}\right],$ (SM31)